Allison:

A tower is composed of a prism with a square base and a pyramid. The base length is 20 meters, and the height of the prism is 40 meters, while the slant height of the pyramid is 10√2 meters. What is the total surface area, including the bottom base?

Allison:

@Bob @rootbeer003 @Jaynator494

Allison:

@EvL

Allison:

@Korianna

joshiscool:

Ah i see still mad

Allison:

@EvL do you know?

Allison:

;-;

Allison:

A. 400√2 + 3,600 m^2 B. 200√2 + 3,600 m^2 C. 400√2 + 400 m^2 D. 4000√2 m^2

Allison:

@EvL Is it A

Allison:

EvL is da best v.v

EvL:

Find the area of the base of the tower first: Since it's a square it will be: \(\sf length \times length = 20 \times 20 \) Now, find the area on the sides of the tower: one side is :\(\sf length \times height= 20 \times 40\) Since you have four sides, you multiply that by 4 Once you find those areas, you add them PLUS the surface area of the square pyramid. To find the area of the sides of the square pyramid : \(\sf one\ side\ area: \frac{1}{2}{b}{h}=\frac{1}{2}(20)(10 \sqrt{2}) \) then you multiply this by 4 since you have FOUR sides of a square pyramid.